### v1

@@ Abstract
Transmission spectroscopy has enabled unprecedented insights into the makeup of exoplanet atmospheres. A transmission spectrum combines contributions from a planet’s morning and evening limbs, but these limbs may have different temperatures, compositions, and aerosol properties due to atmospheric circulation. High-resolution ground-based observations have detected limb asymmetry on several ultra-hot (2000 K) exoplanets, but space-based investigation into limb asymmetry is in its infancy, and limb asymmetry’s prevalence in the broader exoplanet population remains unexplored. We find evidence for limb asymmetry on the exoplanet WASP-107b via transmission spectroscopy from 2.5 to 4.0 micrometers with JWST/NIRCam. This is one of the first low-resolution space-based measurements of limb asymmetry and is unique because, at 770 K, WASP-107b is in a relatively cool regime where planetary terminators are expected to be homogeneous. These observations imply a difference in temperature and cloud properties between WASP-107b’s limbs, challenging our models of limb asymmetry in this cooler regime.
@@ 1.1 Observations
We observed one transit of WASP-107b on January 14th, 2023 using JWST/NIRCam as part of the MANATEE NIRCam GTO program (Obs.8, JWST-GTO-1185, ref.[15]). We simultaneously used NIRCam’s F210M filter in the short-wavelength channel and the F322W2 grism in the long-wavelength channel [16, 17]. For both, this observation was a continuous time-series of 1,293 integrations at 20.22-second cadence using the BRIGHT2 readout pattern on the SUBGRISM256 subarray. We reduced these data using three independent pipelines: Eureka! [18], tshirt [19], and Pegasus [Beatty, T., et al. in review]. Each pipeline produced consistent results, discussed in the Methods, validating our results against systematic biases from the particular reduction method used. We focus primarily on the results from tshirt due to its more effective 1/f noise correction [20]. We produced a band-integrated light curve at 2.093 0.102 m from F210M, as well as a band-integrated (2.45 – 3.95 m) light curve and thirty (R64) spectroscopic light curves from F322W2. Our F322W2 dynamic spectrum and broad-band light curve are shown in Figure 1.
We also collected new and archival transit and radial velocity observations of WASP-107b using other instruments over a wide wavelength range. Uncertainty in the planet’s orbital parameters, particularly the time of conjunction, can bias the signal of limb asymmetry [7, 21, 22]. To combat this, we use these ancillary data to precisely measure WASP-107b’s orbit before our analysis of the JWST/NIRCam data. These include radial velocity measurements with the CORALIE and Keck I/HIRES spectrographs collected by refs.[4, 5]. Also, we downloaded public data covering three transits with the Transiting Exoplanet Survey Satellite, and one transit at 4.5 m with the Infrared Array Camera on the Spitzer Space Telescope (Program 13052, PI: M. Werner). We observed one new transit in the SDSS i-band using the Goodman Spectrograph [23] on the Southern Astrophysical Research Telescope. Lastly, we obtained the broad-band (5 – 12 m) light curve of one transit using JWST’s Mid-Infrared Instrument in Low-Resolution Spectroscopy mode (MIRI/LRS), observed in JWST program 1280 and given to us by PI P.O. Lagage. We describe the reduction and fitting of these data in the Methods, with particular focus on our derivation of WASP-107b’s time of conjunction in Section 3.5. The results are tabulated in Table 1. With these data, we measured WASP-107b’s time of conjunction to a 1- precision of 0.7 seconds.
@@ 1.2.1 Limb transmission spectra
We fit the spectroscopic light curves with asymmetric-limb transit models using the modeling package catwoman [22, 24]. Described further in Methods Section 3.4.4, we fit for the planet’s evening and morning limb planet-star radius ratios at each wavelength, and derived transmission spectra for each limb. We assume the planet’s rotation axis is aligned to its orbit, discussed further in Methods Section 3.4.5. We left the time of conjunction () as a free parameter with a Bayesian prior set to its posterior distribution from fitting our ancillary observations, and did not vary from this prior. The best-fit spectra are shown in Figure 2 and tabulated in Table 2.
@@
As discussed by ref.[9], limb asymmetry can be more confidently attributed to an atmospheric signal, rather than systematics or ephemeris error, if larger asymmetry is observed at wavelengths corresponding to expected features in the transmission spectrum. The larger asymmetry we see around 2.7 m matches an expected water absorption band, which was detected in WASP-107b’s atmosphere by previous HST (6.5 significance) [25] and JWST/MIRI observations (12.5 significance) [26]. The presence of other molecules in WASP-107b’s atmosphere is being explored in a forthcoming paper [Welbanks, L., et al. in review]. There is no reason to suggest water would not be present at both terminators, so the relatively flatter morning limb spectrum is likely the result of that limb being cooler which reduces the size of transmission features. As discussed further later, cloudiness likely plays a role too.
@@ 2 Discussion
To date, limb asymmetry has been detected via ground-based high-resolution transmission spectroscopy on the T 2150 K WASP-76b [33, 34, 35, 36], the 2260 K MASCARA-2b [37], and the 2720 K WASP-121b [38, 39] via molecular abundance and wind speed gradients. WASP-107b joins the 1200 K WASP-39b [Espinoza, N., et al. in review, Delisle, S., et al. in review] as the only planets on which limb asymmetry has been detected via space-based low-resolution transmission spectroscopy, and separate limb transmission spectra recovered. Our discovery on WASP-107b is uniquely surprising because, at an equilibrium temperature of only 770 K, WASP-107b is significantly cooler than these other planets. Such cool atmospheres are not expected to exhibit strong spatial heterogeneity, as their circulation is expected to efficiently distribute heat globally, leading to generally low day-night contrast and homogeneous terminators [e.g. 7, 6, 9]. Our result suggests that either limb asymmetry is more predominant than current models predict, or WASP-107b is an outlier.
@@
We compared our observations to atmospheric models for the morning and evening terminator generated using ScCHIMERA. Described further in Methods Section 3.6, we fixed each model to a metallicity of 10 times the Solar value and carbon-to-oxygen ratio of 0.35. These are the bestfit values from a combined retrieval analysis of previous HST/WFC3 and JWST/MIRI observations [26], and are very consistent with the additional retrievals of Welbanks, et al. (in review). We include a vertically uniform gray “cloud”. The clouds or hazes in WASP-107b’s atmosphere are likely more complex than this, but this simple treatment enables initially probing the relative role of clouds/hazes without adding more free parameters than is justified. We fit a grid of temperature and cloud opacity to the evening and morning limb spectra individually using the -minimization method [40, 41].
@@ 3.2.2 Reduction with Eureka!
Our reduction of the NIRCam F322W2 data with Eureka! followed the same procedure as described by ref.[44], with the only differences being package version numbers. In particular, we used version 0.10.dev15+g80126b56.d20230613 of Eureka![18], CRDS version 11.17.0 and context 1093, and jwst package version 1.10.2. The exact Eureka! Control Files and Eureka! Parameter Files we used are available for download (Zenodo link to be created on final submission). In summary, we cropped the frame to focus on relevant pixels, corrected for the curvature of the spectral trace, performed background subtraction per-column and per-integration, and then performed optimal spectral extraction using the pixels within 9 pixels of the flattened spectral trace. We then sigma-clipped the observations compared to a boxcar smoothed version of the observations to remove cosmic rays without removing sharp astrophysical features like the transit ingress/egress.
@@ 3.2.3 Reduction with Pegasus
We also reduced and extracted the transmission spectrum of WASP-107b using the Pegasus pipeline (https://github.com/TGBeatty/PegasusProject). This pipeline will be described in more detail in a forthcoming paper [Beatty, T., et al. in review], and we briefly summarize it here. We began with the rateint files from the jwst pipeline v1.10.2, using CRDS version 11.17.0 and context 1093.
@@ 3.2.4 Comparing Reduction Pipelines
As in our primary analysis, we fit the spectroscopic light curves from each reduction with an asymmetric-limb transit model (see Section 3.4.4 for details) and derived morning and evening limb transmission spectra for each. All three reductions lead to the same qualitative result: an evening-limb spectrum that is generally at higher transit depth and displays stronger spectral features compared to the relatively flat morning-limb spectrum (Extended Data Figure 6). Comparing the transit depths in each individual channel, the morning and evening limb depths are all consistent within 1 between reductions in each individual channel, except for one outlier at 3.77 m where Eureka! and Pegasus differ by 1.08- (Extended Data Figure 6). There are no channels where tshirt disagrees with the other two reductions by more than 1-. The agreement between the evening limb spectra are particularly good, where the depths differ by less than 0.5 mutual standard deviations in most channels. The only notable difference between the three reductions’ results is that the morning limb spectrum derived from tshirt is generally slightly lower than that of PEGASUS and Eureka! at short wavelengths, though never by more than 1- as mentioned. A potential reason for this, as well as our justification of choosing tshirt as our primary focus, is described further below.
@@
We ultimately chose to use tshirt in our primary analysis because, of the three pipelines, t-shirt performs the most robust correction for 1/f noise. This 1/f noise is one of, if not the most, dominant systematic noise source in NIRCam time series data using GRISMR readout [20, 45]. Eureka! and PEGASUS both adopt very similar techniques for 1/f noise correction, including a “reference pixel step” where the mean value of each detector amplifier’s reference (i.e., unexposed) pixels are subtracted from the value of the exposed pixels. This step is meant to eliminate overall bias offsets which may be different between amplifiers, as well as odd/even column effects, pre-amplifier resets, and 1/f noise. However, these reference pixels exhibit noise and do not sample frequencies greater than 200 Hz, reducing the effectiveness of this correction when used alone. Further, this step can occasionally fail, leaving systematic biases including 1/f noise present in the data. In these failure cases, PEGASUS then uses a custom routine hxrg-ref-pixel (linked in Section 3.2.3) to re-correct the data, while Eureka! does no re-correction. On the other hand, t-shirt replaces this reference pixel step with row-by-row, odd/even by amplifier (ROEBA) correction, as described above and in ref.[45]. ROEBA correction better corrects for and reduces 1/f noise in the data than the default correction [20, 45]. This difference in 1/f noise correction is most likely the reason for the slight differences between the spectra derived from tshirt and the other two pipelines, particularly at the shortest wavelengths of the morning limb spectra (Figure 6). A similar slight offset in absolute transit depth between tshirt and Eureka! has been seen in other analyses, and was directly traceable to this 1/f noise correction [44].
@@ 3.3 Ancillary Data Reduction
Finally, JWST/MIRI observed one transit of WASP-107b in Low Resolution Spectroscopy (LRS) mode during program JWST-GTO-1280 (PI: P.O. Lagage). We were given use of the broad-band (5–12 m) light curve of this transit observation for our analysis. The data was reduced using Eureka!, following the same procedure described in the Eureka! v1 reduction section of ref.[52]. We used a different custom linearity correction than ref.[26], but verified that this did not ultimately change our result.
@@ 3.4.1 Limb Darkening Treatment
Before fitting any of the data, we computed the expected limb darkening coefficients for each instrument’s bandpass to use as priors. We used the online Exoplanet Characterization Toolkit Limb Darkening Calculator [53] with the Kurucz ATLAS9 stellar model grid, with parameters for WASP-107 of Teff = 4425 K, log = 4.63 dex, and [Fe/H] = 0.02 dex. We computed the corresponding model quadratic limb darkening coefficients for each instrument’s bandpass. The coefficient values pertaining to the ancillary data fit priors are listed in Table 1, and all values including those used for the spectroscopic F322W2 light curves are listed in Table 3 of the Extended Data. When fit freely with no connection between wavelengths, limb darkening coefficient profiles can take unphysical, non-smooth, non-self-consistent shapes as a function of wavelength. We use these model coefficients as Gaussian priors in our fit of the ancillary data to ensure that the coefficients are physical and self-consistent, while still allowing their uncertainty to be reflected in the ultimate result. We fix these coefficients when fitting the spectroscopic F322W2 data to reduce the number of free parameters, but verified that this decision did not affect our result (Section 1.2.2).

### v2

@@ Abstract
The atmospheric properties of hot exoplanets are expected to be different between the morning and the evening limb due to global atmospheric circulation. Ground-based observations at high spectral resolution have detected this limb asymmetry in several ultra-hot (2000 K) exoplanets, but the prevalence of the phenomenon in the broader exoplanetary population remains unexplored. Here we use JWST/NIRCam transmission spectra between 2.5 and 4.0 µm to find evidence of limb asymmetry on exoplanet WASP-107 b. With its equilibrium temperature of 770 K and low density of 0.126 gm c-3, WASP-107 b probes a very different regime compared to ultra-hot giant planets and was not expected to exhibit substantial spatial heterogeneity according to atmospheric models. We infer instead a morning-evening temperature difference on the order of 100 K with a hotter evening limb. Further observations on other cooler exoplanets are needed to determine whether WASP-107 b is an outlier or the models underestimate the presence of limb asymmetry in exoplanets.
Main Text
@@ 1.1 Observations
We observed one transit of WASP-107b on January 14th, 2023 using JWST/NIRCam as part of the MANATEE NIRCam GTO program (Obs.8, JWST-GTO-1185, ref.[15]). We simultaneously used NIRCam’s F210M filter in the short-wavelength channel and the F322W2 grism in the long-wavelength channel [16, 17]. For both, this observation was a continuous time-series of 1,293 integrations at 20.22-second cadence using the BRIGHT2 readout pattern on the SUBGRISM256 subarray. We reduced these data using three independent pipelines: Eureka! [18], tshirt (https://github.com/eas342/tshirt), and Pegasus [19]. Each pipeline produced consistent results, discussed in the Methods, validating our results against systematic biases from the particular reduction method used. We focus primarily on the results from tshirt due to its more effective 1/f noise correction [20]. We produced a band-integrated light curve at 2.093 0.102 m from F210M, as well as a band-integrated (2.45 – 3.95 m) light curve and thirty (R64) spectroscopic light curves from F322W2. Our F322W2 dynamic spectrum and broad-band light curve are shown in Figure 1.
We also collected new and archival transit and radial velocity observations of WASP-107b using other instruments over a wide wavelength range. Uncertainty in the planet’s orbital parameters, particularly the time of conjunction, can bias the signal of limb asymmetry [7, 21, 22]. To combat this, we use these ancillary data to precisely measure WASP-107b’s orbit before our analysis of the JWST/NIRCam data. These include radial velocity measurements with the CORALIE and Keck I/HIRES spectrographs collected by refs.[4, 5]. Also, we downloaded public data covering three transits with the Transiting Exoplanet Survey Satellite, and one transit at 4.5 m with the Infrared Array Camera on the Spitzer Space Telescope (Program 13052, PI: M. Werner). We observed one new transit in the SDSS i-band using the Goodman Spectrograph [23] on the Southern Astrophysical Research Telescope. Lastly, we obtained the broad-band (5 – 12 m) light curve of one transit using JWST’s Mid-Infrared Instrument in Low-Resolution Spectroscopy mode (MIRI/LRS), observed in JWST program 1280 and given to us by PI P.O. Lagage. We describe the reduction and fitting of these data in the Methods, with particular focus on our derivation of WASP-107b’s time of conjunction in Section 3.5. The results are tabulated in Extended Data Table 1. With these data, we measured WASP-107b’s time of conjunction to a 1- precision of 0.7 seconds.
@@ 1.2.1 Limb transmission spectra
We fit the spectroscopic light curves with asymmetric-limb transit models using the modeling package catwoman [22, 24]. Described further in Methods Section 3.4.4, we fit for the planet’s evening and morning limb planet-star radius ratios at each wavelength, and derived transmission spectra for each limb. We assume the planet’s rotation axis is aligned to its orbit, discussed further in Methods Section 3.4.5. We left the time of conjunction () as a free parameter with a Bayesian prior set to its posterior distribution from fitting our ancillary observations, and did not vary from this prior. The best-fit spectra are shown in Figure 2 and tabulated in Extended Data Table 2.
@@
As discussed by ref.[9], limb asymmetry can be more confidently attributed to an atmospheric signal, rather than systematics or ephemeris error, if larger asymmetry is observed at wavelengths corresponding to expected features in the transmission spectrum. The larger asymmetry we see around 2.7 m matches an expected water absorption band, which was detected in WASP-107b’s atmosphere by previous HST (6.5 significance) [25] and JWST/MIRI observations (12.5 significance) [26]. The presence of other molecules in WASP-107b’s atmosphere is explored in ref.[27]. There is no reason to suggest water would not be present at both terminators, so the relatively flatter morning limb spectrum is likely the result of that limb being cooler which reduces the size of transmission features. As discussed further later, cloudiness likely plays a role too.
@@ 2 Discussion
To date, limb asymmetry has been detected via ground-based high-resolution transmission spectroscopy on the T 2150 K WASP-76b [34, 35, 36, 37], the 2260 K MASCARA-2b [38], and the 2720 K WASP-121b [39, 40] via molecular abundance and wind speed gradients. WASP-107b joins the 1200 K WASP-39b [[41], Delisle, S., et al. in review] as the only planets on which limb asymmetry has been detected via space-based low-resolution transmission spectroscopy, and separate limb transmission spectra recovered. Our discovery on WASP-107b is uniquely surprising because, at an equilibrium temperature of only 770 K, WASP-107b is significantly cooler than these other planets. Such cool atmospheres are not expected to exhibit strong spatial heterogeneity, as their circulation is expected to efficiently distribute heat globally, leading to generally low day-night contrast and homogeneous terminators [e.g. 7, 6, 9]. Our result suggests that either limb asymmetry is more predominant than current models predict, or WASP-107b is an outlier.
@@
We compared our observations to atmospheric models for the morning and evening terminator generated using ScCHIMERA. Described further in Methods Section 3.6, we fixed each model to a metallicity of 10 times the Solar value and carbon-to-oxygen ratio of 0.35. These are the bestfit values from a combined retrieval analysis of previous HST/WFC3 and JWST/MIRI observations [26], and are very consistent with the additional retrievals of ref.[27]. We include a vertically uniform gray “cloud”. The clouds or hazes in WASP-107b’s atmosphere are likely more complex than this, but this simple treatment enables initially probing the relative role of clouds/hazes without adding more free parameters than is justified. We fit a grid of temperature and cloud opacity to the evening and morning limb spectra individually using the -minimization method [42, 43].
@@ 3.2.1 Reduction with tshirt
We fit the resulting lightcurves produced by tshirt in the same method as for all reductions, which is described in Section 3.4.4. The best-fit transmission spectra are compared to those of our other reductions in Supplementary Figure 1.
@@ 3.2.2 Reduction with Eureka!
Our reduction of the NIRCam F322W2 data with Eureka! (https://eurekadocs.readthedocs.io/en/latest/) followed the same procedure as ref.[46], which we summarize here. We used version 0.10.dev15+g80126b56.d20230613 of Eureka![18], CRDS version 11.17.0 and context 1093, and jwst package version 1.10.2. The exact Eureka! Control Files and Eureka! Parameter Files we used are available for download (https://zenodo.org/doi/10.5281/zenodo.12765422).
Eureka!’s Stage 1 and 2 use the jwst pipeline for basic calibration. We ran both stages using their default settings, except for increasing the Stage 1 jump step’s rejection threshold to 6.0 and skipping the photom step in Stage 2. In Stage 3, we cropped the frame to focus on relevant pixels (y-pixels 5–64 and x-pixels 15–1709), corrected for the curvature of the spectral trace, performed background subtraction per-column and per-integration, and then performed optimal spectral extraction using the pixels within 9 pixels of the flattened spectral trace. We then sigma-clipped (10 threshold) the observations compared to a 50-integration wide boxcar smoothed version of the observations to remove cosmic rays without removing sharp astrophysical features like the transit ingress/egress.
We fit the resulting lightcurves produced by Eureka! in the same method as for all reductions, which is described in Section 3.4.4. The best-fit transmission spectra are compared to those of our other reductions in Supplementary Figure 1.
@@ 3.2.3 Reduction with Pegasus
We also reduced and extracted the transmission spectrum of WASP-107b using the Pegasus pipeline (https://github.com/TGBeatty/PegasusProject). This pipeline is described in more detail in ref.[19], and we briefly summarize it here. We began with the rateint files from the jwst pipeline v1.10.2, using CRDS version 11.17.0 and context 1093.
@@
We fit the resulting lightcurves produced by Pegasus in the same method as for all reductions, which is described in Section 3.4.4. The best-fit transmission spectra are compared to those of our other reductions in Supplementary Figure 1.
@@ 3.2.4 Comparing Reduction Pipelines
As in our primary analysis, we fit the spectroscopic light curves from each reduction with an asymmetric-limb transit model (see Section 3.4.4 for details) and derived morning and evening limb transmission spectra for each. All three reductions lead to the same qualitative result: an evening-limb spectrum that is generally at higher transit depth and displays stronger spectral features compared to the relatively flat morning-limb spectrum (Supplementary Figure 1). Comparing the transit depths in each individual channel, the morning and evening limb depths are all consistent within 1 between reductions in each individual channel, except for one outlier at 3.77 m where Eureka! and Pegasus differ by 1.08- (Supplementary Figure 1). There are no channels where tshirt disagrees with the other two reductions by more than 1-. The agreement between the evening limb spectra are particularly good, where the depths differ by less than 0.5 mutual standard deviations in most channels. The only notable difference between the three reductions’ results is that the morning limb spectrum derived from tshirt is generally slightly lower than that of PEGASUS and Eureka! at short wavelengths, though never by more than 1- as mentioned. A potential reason for this, as well as our justification of choosing tshirt as our primary focus, is described further below.
@@
We ultimately chose to use tshirt in our primary analysis because, of the three pipelines, t-shirt performs the most robust correction for 1/f noise. This 1/f noise is one of, if not the most, dominant systematic noise source in NIRCam time series data using GRISMR readout [20, 47]. Eureka! and PEGASUS both adopt very similar techniques for 1/f noise correction, including a “reference pixel step” where the mean value of each detector amplifier’s reference (i.e., unexposed) pixels are subtracted from the value of the exposed pixels. This step is meant to eliminate overall bias offsets which may be different between amplifiers, as well as odd/even column effects, pre-amplifier resets, and 1/f noise. However, these reference pixels exhibit noise and do not sample frequencies greater than 200 Hz, reducing the effectiveness of this correction when used alone. Further, this step can occasionally fail, leaving systematic biases including 1/f noise present in the data. In these failure cases, PEGASUS then uses a custom routine hxrg-ref-pixel (linked in Section 3.2.3) to re-correct the data, while Eureka! does no re-correction. On the other hand, t-shirt replaces this reference pixel step with row-by-row, odd/even by amplifier (ROEBA) correction, as described above and in ref.[47]. ROEBA correction better corrects for and reduces 1/f noise in the data than the default correction [20, 47]. This difference in 1/f noise correction is most likely the reason for the slight differences between the spectra derived from tshirt and the other two pipelines, particularly at the shortest wavelengths of the morning limb spectra (Supplementary Figure 1). A similar slight offset in absolute transit depth between tshirt and Eureka! has been seen in other analyses, and was directly traceable to this 1/f noise correction [46].
@@ 3.3 Ancillary Data Reduction
Finally, JWST/MIRI observed one transit of WASP-107b in Low Resolution Spectroscopy (LRS) mode during program JWST-GTO-1280 (PI: P.O. Lagage). We were granted use of the broad-band (5–12 m) light curve of this transit observation for our analysis. The data was reduced using Eureka!, following the same procedure described in detail in the Eureka! v1 reduction section of ref.[54], which we briefly summarize here. We again use the jwst pipeline for Stage 1 and 2 with its default parameters, except for Stage 1’s jump step rejection threshold which we increased to 8.0, and skipping Stage 2’s photom step. In Stage 3, we extract pixels 11–61 along the spatial direction and 140–393 along the dispersion direction, not including any pixels marked as “DO_NOT_USE” in the data quality array. We then perform background subtraction accounting for MIRI/LRS’ “cruciform artifact” by first removing bright scattered light rays via sigma-clipping, and excess periodic background noise and 1/ noise by doing this subtraction per-column and per-integration. To further account for potential linear trends in the background flux along the detector, as in ref.[54], we used the mean from an equal number of pixels on either side of the spectral trace for each column and integration when computing the background level. Finally, we perform optimal extraction on the spectrum by calculating a median frame, clipping 5 outliers along the time axis, and smoothing the spectrum using a 7-pixel wide boxcar filter. In our reduction, we used a different custom linearity correction than ref.[26] developed for this same data, but we verified that this did not ultimately change our result.
@@ 3.4.1 Limb Darkening Treatment
Before fitting any of the data, we computed the expected limb darkening coefficients for each instrument’s bandpass to use as priors. We used the online Exoplanet Characterization Toolkit Limb Darkening Calculator [55] with the Kurucz ATLAS9 stellar model grid, with parameters for WASP-107 of Teff = 4425 K, log = 4.63 dex, and [Fe/H] = 0.02 dex. We computed the corresponding model quadratic limb darkening coefficients for each instrument’s bandpass. The coefficient values pertaining to the ancillary data fit priors are listed in Extended Data Table 1, and all values including those used for the spectroscopic F322W2 light curves are listed in Supplementary Table 1. When fit freely with no connection between wavelengths, limb darkening coefficient profiles can take unphysical, non-smooth, non-self-consistent shapes as a function of wavelength. We use these model coefficients as Gaussian priors in our fit of the ancillary data to ensure that the coefficients are physical and self-consistent, while still allowing their uncertainty to be reflected in the ultimate result. We fix these coefficients when fitting the spectroscopic F322W2 data to reduce the number of free parameters, but verified that this decision did not affect our result (Section 1.2.2).